# Conformal Tradeoffs: Guarantees Beyond Coverage

### v1

@@ 6 Discussion
This work uses conformal prediction as deployed decision infrastructure, where the object of interest is not a one-off marginal guarantee but the realized behavior of a fixed rule over a finite operational window. In that regime, the key quantities are operational rate vectors—commitment, abstention, and error exposure—that govern downstream workload and harm. The central message is that coverage is special: because it depends only on the rank of the future true-label score among calibration scores (David and Nagaraja, 2003), its calibration-conditional variability admits an exact, distribution-free finite-sample characterization. Most other deployment KPIs do not share this pivot structure; they depend on multi-score geometry and threshold interactions, so validity must be handled by explicit auditing and uncertainty propagation.
The first contribution, Small-Sample Beta Correction (SSBC), addresses a practical mismatch between a user’s requested coverage level and what nominal split conformal delivers when the calibration set is small. The simulations show that nominal grid selection can yield large calibration-conditional violation probabilities, driven purely by calibration randomness, even when exchangeability holds. SSBC restores an operational semantics by selecting the calibration index through inversion of the exact Beta or Beta–Binomial laws, depending on whether one targets the long-run or finite-window coverage event. In contrast, DKWM-style corrections (Massart, 1990) achieve control primarily through conservatism, often pushing the system to overly abstain. SSBC therefore offers a useful middle ground: it is still distribution-free, but it allocates conservatism only where the finite-sample semantics demand it.
The second contribution, Calibrate–and–Audit, makes explicit a fact that is often implicit in practice: once calibrated, a conformal rule induces a finite partition of score space with a small set of region labels, and downstream systems inevitably map those region labels to actions. By estimating the region masses and region–label composition, the audit table marginalizes the calibration-induced geometry into a reusable empirical interface. This enables forward-looking predictive envelopes for operational rates over future windows, and supports scenario planning over attainable trade-offs across calibration settings. The results illustrate that operational fronts can be explored without retraining the base model, and that uncertainty in these fronts is quantifiable when one treats calibration as a one-time random event and deployment as finite-window sampling.
A key conceptual implication is the feasibility coupling induced by a fixed partition. Because the same conformal interface must serve multiple downstream objectives, not all combinations of commit and error exposure are simultaneously achievable by tuning a single scalar calibration parameter. The regime structure observed in trade-off curves is therefore not an artifact of a particular synthetic model; it follows from the constrained geometry of the induced regions. This perspective also clarifies when cost-aware operation is coherent. Action coherence depends on within-region label frequencies; if a region is label-mixed, forced commitment can be cost-incoherent across plausible cost ratios.
@@
Several limitations delineate the scope of these guarantees. First, the distribution-free statements rely on exchangeability; dataset shift or model drift (Tibshirani et al., 2019; Fannjiang et al., 2022) breaks the exact finite-sample laws and must be handled by monitoring, re-calibration (Vovk, 2015; Barber et al., 2021), or domain-adaptive variants. Second, SSBC is tied to the discrete split-conformal grid; unavoidable granularity appears at small $n_{\mathrm{cal}}$, and very strict $(\alpha^{\star},\delta)$ targets may be unattainable without larger calibration sets. Third, the audit-based envelopes inherit variance from region sparsity; rare regions may require pooling, hierarchical smoothing, or targeted data collection. Finally, we focused on binary classification for clarity. Extending the action-coherence analysis to multi-class, structured outputs, or complex abstention policies is an important next step.
Overall, the paper argues for separating coverage guarantees from operational planning, i.e. what workload and harm rates a deployed conformal interface will induce. SSBC supplies the former, while Calibrate–and–Audit supplies the latter. Together, they support a deployment workflow in which coverage is treated as a calibrated contract and the remaining KPIs are treated as audited, forecasted quantities with explicit, distribution-free finite-sample uncertainty.
@@ 7 Conclusion
We studied conformal prediction in the setting that matters for deployment: a single calibrated rule reused over finite operational windows. In this regime, coverage alone admits an exact, distribution-free calibration-conditional law, while other operational KPIs require explicit auditing.
SSBC provides a deterministic calibration-index selection rule that aligns a user’s coverage request with a PAC-style operational semantics under small calibration sets, including the finite-window case via the induced Beta–Binomial law. Calibrate–and–Audit complements this by turning the calibrated conformal predictor into a reusable region–label interface, enabling predictive envelopes for future commitment, abstention, and error-exposure rates and supporting scenario planning over attainable trade-offs.
Taken together, these tools move conformal prediction (Angelopoulos and Bates, 2023) from a single coverage guarantee toward a practical planning framework: coverage is treated as a contract with calibrated semantics, and the rest of the deployed behavior is forecasted and stress-tested with finite-sample uncertainty. Future work should focus on robust monitoring under shift, partition refinement to improve decision coherence, and extensions to richer output spaces and downstream policies.
@@ C.1 Setup and conformal grid
C.2 Exact law of realized (calibration-conditional) coverage
@@ C.3 SSBC objective: calibration-conditional PAC semantics
Selection rule (least conservative admissible grid point).
@@ C.5 Infinite-window limit
Proposition 2 (Infinite-window limit of SSBC).
Fix $n_{\mathrm{cal}}$, $\alpha^{\star}$, and $\delta$. Let $u_{m}$ be the SSBC-selected index obtained by inverting the Beta–Binomial tail for window size $m$, and let $u_{\infty}$ be the SSBC-selected index obtained by inverting the Beta tail. Then
[…]
Proof sketch. Conditional on $p_{\mathrm{cov}}=p$, $\widehat{C}_{m}\to p$ almost surely as $m\to\infty$. Therefore the Beta–Binomial mixture converges weakly to the Beta law for $p_{\mathrm{cov}}$, and the corresponding admissibility conditions for grid points converge. $\square$
@@ C.9 Extended simulation table for calibration-conditional violations
Table 7: Calibration-conditional coverage violation rates with theory. Target miscoverage $\alpha^{\star}=0.10$, confidence $\delta=0.10$. $\alpha_{\mathrm{grid}}$ is the grid point selected on the conformal grid, $\alpha_{\mathrm{cont}}$ is the requested value under the DKWM correction, and $m_{\mathrm{cal}}$ is the calibration-window size. Results are based on $10^{6}$ calibration draws and a finite deployment window of size $m_{\mathrm{infer}}=100$. The Obs column reports $\mathbb{P}(\widehat{C}_{m}<1-\alpha^{\star})$. BetaTheory reports $\mathbb{P}(p_{\mathrm{cov}}<1-\alpha^{\star})$ under $p_{\mathrm{cov}}\sim\mathrm{Beta}(k,u)$. BBTheory reports $\mathbb{P}(\widehat{C}_{m}<1-\alpha^{\star})$ under the induced Beta–Binomial law.
$n_{\mathrm{cal}}$ Method $u$ $\alpha_{\mathrm{grid}}$ Obs BetaTheory BBTheory $\alpha_{\mathrm{cont}}$
@@ Appendix D Single-sample structural coupling, leave-one-out decoupling, and envelope inflation
In practice, an independent audit window is often unavailable. When the same sample is reused both to select the threshold and to estimate operational rates, threshold selection and evaluation become coupled. This appendix records (i) a minimal structural reason for reuse-induced dependence, (ii) a data-efficient remedy—leave-one-out (LOO) recalibration—that provides effective practical decoupling in our regimes, and (iii) an inflation parameter infl that pessimizes predictive envelopes when residual dependence or regime instability remains.
@@ D.1 Why single-sample reuse introduces dependence
Lemma 3 (Order-statistic coupling under reuse).
Assume $S_{1},\dots,S_{n}$ are exchangeable and continuous, and let $\hat{\tau}=S_{(k)}$. Then, conditional on $\hat{\tau}=t$,
@@
Consequently, $\{I_{i}(t)\}_{i=1}^{n}$ are not conditionally independent given $\hat{\tau}=t$, and for any $i\neq j$,
$\mathrm{Cov}\!\left(I_{i}(t),I_{j}(t)\mid\hat{\tau}=t\right)=\frac{k(k-n)}{n^{2}(n-1)}<0\quad(k<n).$
Proof.
Conditioning on $\hat{\tau}=t$ fixes exactly k1k-1 scores strictly below $t$ and one score equal to $t$, hence $\sum_{i=1}^{n}I_{i}(t)=k$ deterministically. The covariance identity follows from exchangeability and counting. ∎
@@
Operational indicators are functions of the deployed prediction set $C_{\hat{\tau}}(X)=\{y:s(X,y)\leq\hat{\tau}\}$ and therefore inherit reuse-induced dependence. If one naïvely treats reuse-based indicators as i.i.d. Bernoulli under a fixed rule, predictive envelopes can become under-dispersed. Empirically in our studies, the dominant distortion is in dispersion: point estimates often remain close to a two-stage reference, while intervals can be too narrow without decoupling or pessimization.
@@ D.4 Predictive envelope constructions from LOO proxies
Predictive Beta–Binomial envelopes (primary approximation).
@@
Hoeffding-type dominance bound (guardrail).
@@
Single-sample reuse induces structural dependence through order-statistic thresholding (Lemma 3), primarily distorting predictive dispersion. LOO recalibration reduces self-influence and empirically recovers audit-style behavior in our regimes. Predictive envelopes are then constructed from pooled LOO proxies using a Beta–Binomial model, with infl providing a monotone knob for controlled pessimization and a Hoeffding bound serving as a conservative guardrail.
@@ Appendix E Binary conformal partitions: regimes, coupling, and rate primitives
This appendix records the geometric facts used in Section 2, Section 3, and Section 4. We work in the calibration-conditional viewpoint: thresholds are treated as fixed (conditional on the realized calibration draw), and all probabilities are taken with respect to the deployment distribution conditional on $\mathcal{D}_{\mathrm{cal}}$.
E.1 Mondrian split conformal as a four-region partition
Let $\mathcal{Y}=\{0,1\}$ and let $s(x,y)$ be a nonconformity score. Mondrian (class-conditional) split conformal produces thresholds $\tau_{0},\tau_{1}$, and the set-valued output is
@@ E.4 Region–label primitives and rate factorizations
$\mu_{r}(\tau):=\Pr(R_{\tau}(X)=r\mid\mathcal{D}_{\mathrm{cal}})=\sum_{y\in\{0,1\}}p_{r,y}(\tau),\qquad\eta_{r}(\tau):=\Pr(Y=1\mid R_{\tau}(X)=r,\ \mathcal{D}_{\mathrm{cal}})=\frac{p_{r,1}(\tau)}{\mu_{r}(\tau)}\quad(\mu_{r}(\tau)>0).$
@@ F.1 Interface primitives: masses and within-region label frequencies
F.2 Region-measurable action conventions and coherence
Let $\mathcal{A}$ be a finite action set and let $L_{\theta}(a,y)$ be a priced consequence (cost or negative utility), parameterized by $\theta\in\Theta$. A deployed convention is a region-measurable policy
@@ Appendix G Explicit region indicators and projection masks
This appendix instantiates the linear “sum selected cells” formalism from Section 2.4 in the binary toy geometry used in Figure 2 and Figure 3 ($R_{\tau}(x)\xrightarrow{\ \pi\ }C(x)$). Region feasibility and regime facts are recorded in Appendix E; here we keep only the explicit projection masks used in audit computations.
@@ G.3 Binary projection masks
We list four masks used repeatedly in the paper.
@@
(2) Missed positive mass (Figure 1: $q_{10}$ under $\pi_{\mathrm{SI}}$).
In Figure 1, $q_{10}$ is the positive-class mass in region $r_{10}$ (output $\{0\}$ under $\pi_{\mathrm{SI}}$):
$q_{10}(\theta)=\Pr(Y=1,\ C(X)=\{0\})=\Pr(Y=1,\ R_{\tau}(X)=r_{10})=p_{10,1}(\theta),$
with single-cell mask
$G_{10}(\pi_{\mathrm{SI}})=\begin{array}[]{c|cc}&y=0&y=1\\ \hline\cr r_{10}&0&1\\ r_{11}&0&0\\ r_{01}&0&0\\ r_{00}&0&0\end{array}.$
(3) Hedged positive mass (Figure 1: $q_{11}$ under $\pi_{\mathrm{SI}}$).
Similarly,
$q_{11}(\theta)=\Pr(Y=1,\ C(X)=\{0,1\})=\Pr(Y=1,\ R_{\tau}(X)=r_{11})=p_{11,1}(\theta),$
with mask
$G_{11}(\pi_{\mathrm{SI}})=\begin{array}[]{c|cc}&y=0&y=1\\ \hline\cr r_{10}&0&0\\ r_{11}&0&1\\ r_{01}&0&0\\ r_{00}&0&0\end{array}.$
(4) Abstention mass under the commit–reject policy.
Under $\pi_{\mathrm{CR}}$, abstention occurs in regions $r_{11}$ and $r_{00}$, regardless of label. Thus
$G_{\mathrm{abs}}(\pi_{\mathrm{CR}})=\begin{array}[]{c|cc}&y=0&y=1\\ \hline\cr r_{10}&0&0\\ r_{11}&1&1\\ r_{01}&0&0\\ r_{00}&1&1\end{array}$
and
$r_{\mathrm{abs}}(\theta;\pi_{\mathrm{CR}})=\big(p_{11,0}(\theta)+p_{11,1}(\theta)\big)+\big(p_{00,0}(\theta)+p_{00,1}(\theta)\big).$
@@ H.1 Tox21 dataset
The full dataset composition used in this study is reported in the main text (Table 2). We do not duplicate that table here.
H.2 Aggregated coverage summary
The aggregated coverage and prediction-set summary is reported in the main text (Table 3); we do not duplicate that table here.
H.3 Representative endpoint-level operational summaries
@@
H.4 Representation and model protocol

### v2

@@ 6 Discussion
In a deployed conformal system, the object of interest is not only marginal coverage but the realized behavior of a fixed calibrated rule over a finite operational window. In that setting, coverage is special: because it depends only on the rank of the future true-label score among calibration scores, its calibration-conditional variability admits an exact distribution-free finite-sample characterization. Most other deployment quantities, such as commitment, abstention, and error exposure, do not share this pivot structure; they depend on score geometry and threshold interactions, and therefore require explicit auditing and uncertainty quantification.
One consequence is that calibration should be viewed as selecting a fixed interface rather than merely a scalar guarantee. Once thresholds are chosen, the induced conformal partition constrains which operational profiles are attainable. Not all combinations of commitment, deferral, and decisive error exposure can be reached by tuning a single calibration parameter, because changing thresholds reallocates mass across a finite set of region types rather than adjusting each KPI independently. The regime structure seen in the trade-off curves is therefore a property of the induced geometry, not an artifact of a particular synthetic example.
This viewpoint also clarifies the roles of SSBC and Calibrate–and–Audit. SSBC addresses the mismatch between a user’s requested coverage level and what nominal split conformal delivers when calibration sets are small, by selecting the calibration index through inversion of the exact Beta or Beta–Binomial laws. The audit step addresses a different issue: for quantities beyond coverage, uncertainty must be propagated through the realized region–class label table rather than through a rank pivot. In practice, this produces a way to compare calibration settings, quantify finite-window uncertainty for several KPIs at once, and examine cost-coherence of downstream conventions without retraining the base model.
The same geometry also explains why cost-aware operation can fail. Whether a downstream action convention is coherent depends on the within-region label frequencies, not on the conformal set label by itself. A singleton region may still support rejection, or even the opposite commitment, for plausible cost ratios. In that sense, the audit table is not only a device for uncertainty quantification; it also makes explicit which downstream actions are compatible with the information content of the deployed interface.
@@
Several limitations delimit the scope of these results. We focused on binary classification for clarity; extending the same analysis to multi-class outputs, structured predictions, or more complex abstention policies is an important next step. Audit-based predictive envelopes can also become noisy when region masses are sparse, so rare regions may require pooling, hierarchical regularization, or additional data collection. SSBC remains tied to the discrete split-conformal grid, so very strict $(\alpha^{\star},\delta)$ targets may be unattainable at small $n_{\mathrm{cal}}$. More broadly, all of the distribution-free statements rely on exchangeability; dataset shift or model drift (Tibshirani et al., 2019; Fannjiang et al., 2022) breaks the exact finite-sample laws and must be addressed through monitoring, recalibration, or shift-aware variants (Vovk, 2015).
Taken together, these results point to a useful division of labor. Coverage can be calibrated exactly in the split-conformal setting, but broader questions about how the deployed system behaves must be answered by auditing the realized interface. That distinction matters whenever deployment decisions depend on operational trade-offs rather than on coverage alone.
@@ 7 Conclusion
The main message of the paper is that a deployed conformal predictor should be treated as a fixed operational interface, not only as a coverage certificate. In the setting studied here, SSBC calibrates the coverage target at finite sample, while Calibrate–and–Audit is used to evaluate the resulting rule in terms that matter for use: commitment, abstention, error exposure, and decision coherence.
The present analysis is intentionally narrow. It relies on exchangeability and focuses on binary classification, so extending the same operational viewpoint to shift-robust procedures, richer output spaces, and more flexible downstream policies is the natural next step.
Acknowledgements
This work was supported by the U.S. Department of Energy, Office of Science, under Contract No. DE-AC02-05CH11231. Support also came in part from the Berkeley Synchrotron Infrared Structural Biology Program (BSISB), funded by the Biological and Environmental Research Division of the U.S. Department of Energy Office of Science, and from Laboratory Directed Research and Development (LDRD) funding. Additional support came from the U.S. Department of Energy, Office of Science, Scientific Discovery through Advanced Computing (SciDAC) program FORUM-AI for first-principles calculations and AI model development.
@@ C.1 Setup and conformal grid
C.2 Exact law of realized coverage
@@ C.3 SSBC objective: calibration-conditional PAC semantics
Selection rule: least conservative admissible grid point.
@@ C.5 Infinite-window limit
This limiting regime is identified by Marques (2025). In our notation, if $u_{m}$ denotes the SSBC-selected grid index obtained by inverting the Beta–Binomial tail for window size $m$, and $u_{\infty}$ denotes the index obtained from the Beta limit law, then $u_{m}\to u_{\infty}$ as $m\to\infty$. Intuitively, conditional on $p_{\mathrm{cov}}=p$, the empirical coverage $\widehat{C}_{m}$ converges almost surely to $p$, so the finite-window Beta–Binomial criterion approaches the infinite-window Beta criterion.
@@ C.9 Extended simulation table for calibration-conditional violations
Table 7: Calibration-conditional coverage violation rates with theory. Target miscoverage $\alpha^{\star}=0.10$, confidence $\delta=0.10$. $\alpha_{\mathrm{grid}}$ is the grid point selected on the conformal grid, $\alpha_{\mathrm{cont}}$ is the requested value under the DKWM correction, and $m_{\mathrm{cal}}$ is the calibration-window size. Results are based on $10^{6}$ calibration draws and a finite deployment window of size $m_{\mathrm{infer}}=100$. The Obs column reports $\mathbb{P}(\widehat{C}_{m}<1-\alpha^{\star})$. Beta reports $\mathbb{P}(p_{\mathrm{cov}}<1-\alpha^{\star})$ under $p_{\mathrm{cov}}\sim\mathrm{Beta}(k,u)$. BetaBinom reports $\mathbb{P}(\widehat{C}_{m}<1-\alpha^{\star})$ under the induced Beta–Binomial law.
$n_{\mathrm{cal}}$ Method $u$ $\alpha_{\mathrm{grid}}$ Obs Beta BetaBinom $\alpha_{\mathrm{cont}}$
@@ Appendix D Single-sample structural coupling, leave-one-out decoupling, and envelope inflation
In practice, an independent audit set is not always be available. When the same sample is reused both to select the threshold and to estimate operational rates, threshold selection and evaluation become coupled. This appendix records (i) a minimal structural reason for reuse-induced dependence, (ii) a data-efficient remedy—leave-one-out (LOO) recalibration—that provides effective practical decoupling in our regimes, and (iii) an inflation parameter infl that pessimizes predictive envelopes when residual dependence or regime instability remains.
@@ D.1 Why single-sample reuse introduces dependence
Insight from counting.
If $\hat{\tau}=t$ is the $k$th order statistic, then exactly k1k-1 scores lie strictly below $t$ and one score equals $t$ (under continuity, almost surely), so
@@
This already shows that the indicators $\{I_{i}(t)\}_{i=1}^{n}$ cannot be conditionally independent given $\hat{\tau}=t$: once some are known to be one, fewer ones remain available for the others. By exchangeability, for any $i\neq j$,
$\mathrm{Cov}\!\left(I_{i}(t),I_{j}(t)\mid\hat{\tau}=t\right):=\frac{k(k-n)}{n^{2}(n-1)}<0\quad(k<n).$
@@
Operational indicators are functions of the deployed prediction set $C_{\hat{\tau}}(X)=\{y:s(X,y)\leq\hat{\tau}\}$ and therefore inherit reuse-induced dependence. In particular, the nonzero conditional covariance shows that these indicators are not Bernoulli draws under a fixed rule once threshold selection and evaluation are performed on the same sample. To recover the Bernoulli sampling picture needed for the two-stage guarantees in the main text, one therefore needs an independent audit set after calibration. If one naïvely treats reuse-based indicators as i.i.d. Bernoulli under a fixed rule, predictive envelopes can become under-dispersed. Empirically in our studies, the dominant distortion is in dispersion: point estimates often remain close to a two-stage reference, while intervals can be too narrow without decoupling or pessimization.
@@ D.4 Predictive envelope constructions from LOO proxies
Predictive Beta–Binomial envelopes.
@@
Hoeffding-type dominance bound.
@@
Single-sample reuse induces structural dependence through the thresholding mechanism described above, primarily distorting predictive dispersion. LOO recalibration reduces self-influence and empirically recovers audit-style behavior in our regimes. Predictive envelopes are then constructed from pooled LOO proxies using a Beta–Binomial model, with infl providing a monotone knob for controlled pessimization and the Hoeffding bound serving as a conservative guardrail.
@@ Appendix E Binary conformal partitions: regimes, coupling, and rate primitives
This appendix records the geometric facts used in Section 2, Section 3, and Section 4. Here we assume thresholds are fixed and all probabilities are taken with respect to the deployment distribution conditional on $\mathcal{D}_{\mathrm{cal}}$.
E.1 Class-conditional split conformal as a four-region partition.
Let $\mathcal{Y}=\{0,1\}$ and let $s(x,y)$ be a nonconformity score. Class-conditional split conformal produces thresholds $\tau_{0},\tau_{1}$, and the set-valued output is
@@ E.4 Region–label primitives and rate factorizations
$\mu_{r}(\tau)$ $:=\Pr(R_{\tau}(X)=r\mid\mathcal{D}_{\mathrm{cal}})=\sum_{y\in\{0,1\}}p_{r,y}(\tau),$
$\eta_{r}(\tau)$ $:=\Pr(Y=1\mid R_{\tau}(X)=r,\ \mathcal{D}_{\mathrm{cal}})=\frac{p_{r,1}(\tau)}{\mu_{r}(\tau)}\quad(\mu_{r}(\tau)>0).$
@@ F.1 Interface primitives: masses and within-region label frequencies
F.2 Region-associated action conventions and coherence
Let $\mathcal{A}$ be a finite action set and let $L_{\theta}(a,y)$ be a priced consequence (cost or negative utility), parameterized by $\theta\in\Theta$. A deployed convention is a region-associated policy
@@ Appendix G Explicit region indicators and projection masks
This appendix briefly instantiates the linear “sum selected cells” formalism from Section 2.4 in the binary toy geometry used in Figure 2 and Figure 3 ($R_{\tau}(x)\xrightarrow{\ \pi\ }C(x)$). Region feasibility and regime facts are recorded in Appendix E; here we keep only the explicit projection masks needed for audit computations.
@@ G.3 Binary projection masks
We record one detailed worked example and then list several shorter special cases used elsewhere in the paper.
@@
Other projections used in the paper.
The same mask formalism gives the remaining quantities used in Figure 1 and in the binary operational examples:
$\text{missed positive mass:}\qquad q_{10}(\theta)=\Pr(Y=1,\ C(X)=\{0\})=\Pr(Y=1,\ R_{\tau}(X)=r_{10})=p_{10,1}(\theta),$
$\text{hedged positive mass:}\qquad q_{11}(\theta)=\Pr(Y=1,\ C(X)=\{0,1\})=\Pr(Y=1,\ R_{\tau}(X)=r_{11})=p_{11,1}(\theta),$
$\text{abstention mass under }\pi_{\mathrm{CR}}:\qquad r_{\mathrm{abs}}(\theta;\pi_{\mathrm{CR}})=\big(p_{11,0}(\theta)+p_{11,1}(\theta)\big)+\big(p_{00,0}(\theta)+p_{00,1}(\theta)\big).$
These correspond respectively to a one-cell projection, another one-cell projection, and a two-row projection. Their binary masks are obtained immediately by placing ones on the selected cells.
@@ H.1 Tox21 dataset
The full dataset composition and aggregated coverage summary used in this study is reported in the main text (Table 2).
H.2 Representative endpoint-level operational summaries
@@
H.3 Representation and model protocol